# DualPM: Dual Posed-Canonical Point Maps for 3D Shape and Pose Reconstruction

### v3

@@ 3.2 Amodal point maps
A limitation of the dual point maps is that they only capture the visible portion of the object. We now show how to predict a complete 3D point cloud for the object, including the points that are not visible due to self-occlusion.
To this end, we introduce the concept of an amodal point map. Namely, we associate each pixel $\boldsymbol{u}\in M$ within the object mask to the sequence of 3D points $(\boldsymbol{p}_{1},\boldsymbol{p}_{2},\dots)$ that intersect the camera ray through $\boldsymbol{u}$ in order of increasing distance from the camera center. Hence, $\boldsymbol{p}=\boldsymbol{p}_{1}$ is the imaged point considered in Sec. 3.1, $\boldsymbol{p}_{2}$ is the point that would be imaged should we pluck a hole in the object surface at $\boldsymbol{p}_{1}$, and so on. Then, the goal is to predict the amodal point map $\mathcal{P}$, defined as the map that associates to each pixel $\boldsymbol{u}$ the sequence of 3D points $\mathcal{P}(\boldsymbol{u})=(\boldsymbol{p}_{1},\boldsymbol{p}_{2},\dots)$.
Note that the length of the sequence $\mathcal{P}(\boldsymbol{u})$ is not fixed but depends on the pixel $\boldsymbol{u}$. The sequence length is always even because the ray must exit the object once it enters it, as we assume the camera is positioned outside the object. To account for this, we extend the network $\Phi_{P}$ by first predicting a pair of points $(\boldsymbol{p}_{1},\boldsymbol{p}_{2})_{1}$ for each pixel $\boldsymbol{u}$. Subsequently, additional sets of outputs $(\boldsymbol{p}_{2k-1},\boldsymbol{p}_{2k})_{k}$ are added for $k\in[2,K]$, where $K$ denotes the total number of pairs, capturing further potential ray intersections with the object. Unlike the first set of intersections, these additional intersections may not always occur; therefore, we also predict opacities $(\sigma_{2k-1},\sigma_{k})_{k}$ for these cases. As a result, each pixel $\boldsymbol{u}$ is mapped to $3\times 2+(3+1)\times 2\times(K-1)$ scalar outputs. The network $\Phi_{Q}$ is extended in a similar manner. The losses $\mathcal{L}_{P}$ and $\mathcal{L}_{Q}$ are also extended to account for the amodal point maps. We also add additional loss $\mathcal{L}_{\sigma}$ equivalent to those of $\mathcal{L}_{P}$ and $\mathcal{L}_{Q}$ to supervised the opacities prediction.
@@
To efficiently represent the amodal point map for the neural network prediction, we introduce an image-based layered representation for DualPMs denoted as $(P^{*},Q^{*})$. In this representation, the first layer encodes the visible points, and each consecutive layer captures the next set of points occluded by the previous layer. We also represent the opacity maps $\sigma^{*}$ in the same layered fashion.
Since we supervise our method with synthetic data, we can easily generate training targets for this representation. Each vertex of the object mesh is assigned a 6D attribute, consisting of the calculated posed vertex 3D position and the corresponding vertex 3D position in canonical space. The mesh is then rendered from the camera’s viewpoint. During the rendering process, the rasterizer generates a set of pixel-aligned layers, $L_{i}\in\mathbb{R}^{6\times H\times W}$, which contain the rendered 6D attributes, together with the associated opacity map $\hat{\sigma}_{i}^{*}\in[0,1]^{H\times W}$ for every ray-object intersection. We iterate through the rasterized layers from the front to the back until the number of layers specified for our model is reached. This approach yields ground-truth DualPMs with a number of layers corresponding to the ray-object intersections. We split the attributes of each layer corresponding to the canonical and posed positions, yielding ground-truth layered canonical point maps
@@
representations with their associated opacity maps $\sigma_{i}^{*}$. The number of layers $N$ corresponds to $2K$. The canonical $Q$ and posed $P$ point maps can be then simply extracted from this representation by masking out transparent points.
@@ 3.3 Training
Figure 4: Synthetic training data. We generate synthetic training data by rendering a single rigged 3D horse model with two texture variations (top row) in various poses with different environmental maps and under random viewpoints (bottom row).
We train our model using synthetic data generated from a single articulated 3D horse model sourced from the Animodel dataset [17]. Following the data generation pipeline described in Animodel, we randomly sample object poses, camera poses, and illumination conditions. These are modeled by randomly sampling from collection environment maps. For each sample, we render the object into an image $I$ and generate the corresponding image-based layered paired point maps $P^{*}$ and $Q^{*}$, which serve as targets for the model, using the procedure detailed in Sec. 3.2. We illustrate our synthetic training data in Fig. 4.
@@ 4 Experiments
Figure 5: Comparison with the state-of-the-art. We compare our method with the state-of-the-art method 3D-Fauna [32]. For visualization, we paint each point of the recovered point map $P$ with its canonical coordinate. Our reconstructed shape aligns more closely with the object’s structure compared to 3D-Fauna and adapts well to large variations within the object category, despite being trained on a single adult male horse model. Our method reliably recovers shapes outside its training distribution, such as foals and dwarf horses (3rd and 4th rows from the bottom). It also performs well on very challenging poses, such as those shown in the last two rows, where 3D-Fauna completely fails.
@@
Ours 73.2 63.0 64.2 3.13 ± $\pm\ 2.13$ 4.74 ± $\pm\ 1.40$ 2.32 ± $\pm\ 0.78$ 1.11 ± $\pm\ 0.73$ 1.67 ± $\pm\ 0.55$ 2.03 ± $\pm\ 0.71$
Table 1: Quantitative evaluation. We evaluate on PASCAL VOC, reporting PCK@0.1 (higher is better $\uparrow$), and on Animodel [17], reporting the bi-directional Chamfer Distance in centimeters (lower is better $\downarrow$). Our model, trained solely on data from a single horse model, outperforms state-of-the-art approaches trained on each of the respective categories.
We evaluate our DualPM in terms of both qualitative (Sec. 4.1) and quantitative performance (Sec. 4.2) on the horse category, demonstrate zero-shot generalization to unseen categories of quadrupeds (Fig. 6), show that our DualPM can be used to fit 3D skeletons for motion transfer and animations (Sec. 4.4), and conduct an ablation study to assess the impact of our design choices on model performance (Sec. 4.5). We include technical and implementation details in App. A.
@@ 4.1 Qualitative evaluation
We qualitatively evaluate our method on challenging in-the-wild images of horses from the PASCAL VOC dataset [11]. The results are presented in Figs. 1 and 5. Our method successfully recovers the 3D shape and pose of horses from a single image, despite wide variations in pose, shape, and appearance within the horse category, including foals and dwarf horses, even though it is trained on a single adult horse model. We also compare our method with the state-of-the-art approach for single-view reconstruction of deformable objects, 3D-Fauna [32], and present side-by-side comparisons in Fig. 5. Our method performs better in recovering the shape and pose than 3D-Fauna, which struggles to follow the object’s structure accurately and fails to reconstruct intricate details specific to each object.
@@ 4.2 Quantitative evaluation
Following prior work on deformable object reconstruction [25, 29, 71, 32], we evaluate our method on the keypoint transfer task using PASCAL VOC dataset [11]. We adhere to the protocol of [25] and report the percentage of correctly transferred keypoints with a 10% threshold (PCK@0.1) between pairs of images. The results are presented in Tab. 1, where we compare our method with state-of-the-art approaches.
We also measure the bi-directional Chamfer distance between the predicted point maps and the ground truth 3D shape on the Animodel dataset introduced in Farm3D [17] . This dataset comprises renderings of 3D models of animals, including horses, cows, and sheep, in various poses, representing the shapes of both male and female specimens. As our model is trained using renderings of the male horse from this dataset, we evaluate its performance, along with other approaches, on a subset containing only the female horse. To ensure a fair comparison, we also exclude the poses used in this benchmark from our training set. The performance of our method and others is reported in Tab. 1.
Our method consistently outperforms all other approaches on both benchmarks, demonstrating its ability to generalize to unseen poses and shapes. Moreover, it exhibits exceptional generalization on real-world images despite being trained solely on synthetic data.
4.3 Generalization to unseen categories
Figure 6: Results on unseen categories. Our method demonstrates robust generalization to the unseen categories such as cow and sheep, despite being trained solely with a single horse model.
Given the generalization capabilities of our method demonstrated on the horse category, we extend our analysis to unseen quadruped categories, specifically cows and sheep. We evaluate our approach using the same datasets as for the horses, PASCAL VOC [11] and Animodel [17], following the same evaluation protocol, with results reported in Tab. 1. Furthermore, we provide additional qualitative results on the same dataset in Fig. 6. Our method exhibits strong zero-shot generalization to these categories, outperforming state-of-the-art approaches on both datasets, despite being trained exclusively on a single horse model.
@@ 4.4 Skeleton fitting and animation
For each point $(\boldsymbol{p}_{i},\boldsymbol{q}_{i})$ of the DualPM $(P,Q)$, we identify the closest vertex of the mesh in the canonical space to the canonical point $\boldsymbol{q}_{i}$ then we associate the point $(\boldsymbol{p}_{i},\boldsymbol{q}_{i})$ with the skinning weights of that vertex. For each bone $b\in B$, we identify subsets of $\{(\boldsymbol{p}_{i},\boldsymbol{q}_{i})_{b}\}$, where the associated skinning weights between the bone $b$ and the points $(\boldsymbol{p}_{i},\boldsymbol{q}_{i})$ are higher than 0.5. We then compute the transformation that aligns the subset of canonical points $\{(\boldsymbol{q}_{i})_{b}\}$ to the subset of posed points $\{(\boldsymbol{p}_{i})_{b}\}$ using the Procrustes algorithm. This transformation is then applied to the bone $b$ in the canonical space to derive its corresponding pose in the posed space. This process is repeated for all bones in the skeleton to fit the 3D skeleton to the posed points. We show examples of our fitted skeletons in Fig. 1.
@@ 4.5 Ablation study
Additionally, we investigate the effect of conditioning the predictions of the posed point maps $P$ on the input image features $F$ and the canonical point maps $Q$. When the model is conditioned on both the image features $F$ and the canonical point maps $Q$, it tends to overfit to the features of the training synthetic dataset. In contrast, conditioning the posed point maps $P$ solely on the canonical point maps $Q$ leads to better generalization on out-of-distribution images, as we also demonstrate qualitatively in Fig. 8.
@@
Figure 8: Effect of different conditioning schemes. Conditioning solely on the canonical point maps $Q$, as opposed to using both the image features $F$ and the canonical point maps $Q$, results in better generalization. This improvement is particularly evident on extreme out-of-distribution images (b).
@@
Table 2: Ablation study on Animodel Horses. We report the bi-directional Chamfer Distance in centimeters (lower is better $\downarrow$) for Animodel Horses (Female subset). The evaluation follows the same protocol as in Tab. 1.
5 Limitations
Figure 9: Typical failure cases. We illustrate representative failure cases caused by (a) extreme viewpoints, (b) shapes and poses far from the training distribution, and (c-d) inaccuracies in the object segmentation masks.
Despite demonstrating surprising generalization, a current limitation of our method is that any additional synthetic 3D models added to the training dataset would have to be in the same canonical space as the training data. Addressing the challenging problem of aligning the canonical spaces of multiple 3D models would allows us to train on significantly larger datasets which could in turn lead to significant gains in performance for our proposed model. Another limitation of our method is that it is not specifically trained to handle occlusions caused by other objects. This is a limitation shared with other methods, such as 3D-Fauna [71, 32]. To address this, we plan to extend the data generation pipeline to include synthetic occlusions. Additionally, as the 3D reconstruction problem is often ambiguous for the unseen parts of objects, our method predicts only the expectation over all possible reconstructions, which can lead to unrealistic results for the invisible regions. We illustrate our typical failure cases in Fig. 9.
@@ 6 Conclusion
We presented DualPMs, a novel representation for monocular 3D shape and pose reconstruction of deformable objects. Predicting DualPMs is straightforward for a neural network, and they enable solving a variety of geometric tasks for deformable objects, including not only 3D shape and pose reconstruction but also 3D keypoint localization, skeleton fitting and motion transfer. Despite being trained solely on synthetic data generated from a single horse, our method generalizes effectively to real images, including other categories of quadrupeds, and outperforms state-of-the-art methods on the task of 3D shape and pose reconstruction of horses, cows, and sheep.
Acknowledgments.  We thank Paul Engstler helpful discussions. Ben Kaye is supported by the ESPSRC Centre for Doctoral Training in Autonomous Intelligent Machines and Systems, EP/S024050/1. Tomas Jakab and Andrea Vedaldi are supported by ERC-CoG UNION 101001212. Tomas Jakab is also supported by VisualAI EP/T028572/1.
@@
Appendix A Technical details

### v4

@@ 3.2 Amodal point maps
A limitation of dual point maps is that they only capture the visible portion of the object. We now show how to predict a complete 3D point cloud for the object, including points that are not visible due to self-occlusion.
To this end, we introduce the concept of an amodal point map. Namely, we associate each pixel $\boldsymbol{u}\in M$ within the object mask to the sequence of 3D points $(\boldsymbol{p}_{1},\boldsymbol{p}_{2},\dots)$ that intersect the camera ray through $\boldsymbol{u}$ in order of increasing distance from the camera center. Hence, $\boldsymbol{p}=\boldsymbol{p}_{1}$ is the imaged point considered in Sec. 3.1, $\boldsymbol{p}_{2}$ is the point that would be imaged if a hole were plucked in the object surface at $\boldsymbol{p}_{1}$, and so on. The goal is to predict the amodal point map $\mathcal{P}$, defined as the map that associates each pixel $\boldsymbol{u}$ with the sequence of 3D points $\mathcal{P}(\boldsymbol{u})=(\boldsymbol{p}_{1},\boldsymbol{p}_{2},\dots)$.
The length of the sequence $\mathcal{P}(\boldsymbol{u})$ is not fixed but depends on the pixel $\boldsymbol{u}$. The sequence length is always even because the ray must exit the object once it enters it, as we assume the camera is positioned outside the object. To account for this, we extend the network $\Phi_{P}$ to first predict a pair of points $(\boldsymbol{p}_{1},\boldsymbol{p}_{2})_{1}$ for each pixel $\boldsymbol{u}$. Subsequently, additional sets of outputs $(\boldsymbol{p}_{2k-1},\boldsymbol{p}_{2k})_{k}$ are added for $k\in[2,K]$, where $K$ denotes the total number of pairs, capturing further potential ray intersections with the object. Unlike the first set of intersections, these additional intersections may not always occur; therefore, we also predict opacities $(\sigma_{2k-1},\sigma_{k})_{k}$ for these cases. As a result, each pixel $\boldsymbol{u}$ is mapped to $3\times 2+(3+1)\times 2\times(K-1)$ scalar outputs. The network $\Phi_{Q}$ is extended similarly. The losses $\mathcal{L}_{P}$ and $\mathcal{L}_{Q}$ are also extended to account for the amodal point maps. Additionally, we add a loss $\mathcal{L}_{\sigma}$ to supervise the opacity predictions.
@@
To efficiently represent the amodal point map for neural network prediction, we introduce an image-based layered representation for DualPMs, denoted as $(P^{*},Q^{*})$. In this representation, the first layer encodes the visible points, and each consecutive layer captures the next set of points occluded by the previous layer. We also represent the opacity maps $\sigma^{*}$ in the same layered fashion.
Since we supervise our method with synthetic data, we can easily generate training targets for this representation. Each vertex of the object mesh is assigned a 6D attribute, consisting of the posed vertex’s 3D position and the corresponding vertex’s 3D position in canonical space. The mesh is then rendered from the camera’s viewpoint. During the rendering process, the rasterizer generates a set of pixel-aligned layers, $L_{i}\in\mathbb{R}^{6\times H\times W}$, which contain the rendered 6D attributes, together with the associated opacity map $\hat{\sigma}_{i}^{*}\in[0,1]^{H\times W}$ for every ray-object intersection. We iterate through the rasterized layers from the front to the back until the number of layers specified for our model is reached. This approach yields ground-truth DualPMs with a number of layers corresponding to the ray-object intersections. We split the attributes of each layer corresponding to the canonical and posed positions, yielding ground-truth layered canonical point maps
@@
with their associated opacity maps $\sigma_{i}^{*}$. The number of layers $N$ corresponds to $2K$. The canonical $Q$ and posed $P$ point maps can then be extracted from this representation by masking out transparent points.
@@ 3.3 Training
Figure 4: Synthetic training data. We generate synthetic training data by rendering a rigged 3D model of a category in various poses with different environmental maps and under random viewpoints.
We train a separate model for each category using synthetic data generated from one or two articulated 3D assets per category sourced from the Animodel dataset [18]. We use a separate model for each biological sex in cases where animals exhibit distinct anatomical structures specific to males and females, such as horns in cattle. Following the data generation pipeline described in Animodel, we randomly sample object poses, camera poses, and illumination conditions. These are modeled by randomly sampling from a collection of environment maps. For each sample, we render the object into an image $I$ and generate the corresponding image-based layered paired point maps $P^{*}$ and $Q^{*}$, which serve as targets for the model, using the procedure detailed in Sec. 3.2. We illustrate our synthetic training data in Fig. 4.
@@ 4 Experiments
Figure 5: Comparison with the state-of-the-art. We compare our method with 3D-Fauna [33]. For visualization, we paint each point of the recovered point map $P$ with its canonical coordinate. Our reconstructed shape aligns more closely with the object’s structure compared to 3D-Fauna and adapts well to large variations within the object category. Our method reliably recovers shapes outside its training distribution, such as foals and donkeys (rows 3 and 4). It also performs well on very challenging poses (rows 1 and 2), where 3D-Fauna completely fails.
@@
Trellis [73] 5.11 ± $\pm\ 3.53$ 5.50 ± $\pm\ 3.37$ 2.23 ± $\pm\ 1.39$ 1.85 ± $\pm\ 1.31$ 1.93 ± $\pm\ 1.23$ 1.96 ± $\pm\ 1.22$
Ours 73.2 66.86 66.82 3.13 ± $\pm\ 2.13$ 2.63 ± $\pm\ 1.23$ 1.52 ± $\pm\ 0.88$ 1.11 ± $\pm\ 0.73$ 0.92 ± $\pm\ 0.43$ 1.33 ± $\pm\ 0.78$
Table 1: Quantitative evaluation. We evaluate on PASCAL VOC, reporting PCK@0.1 (higher is better $\uparrow$), and on Animodel [18], reporting the bi-directional Chamfer Distance in centimeters (lower is better $\downarrow$). Our model, trained solely on data from one or two models per category, outperforms other state-of-the-art approaches.
We evaluate our DualPM in terms of both qualitative (Sec. 4.1) and quantitative performance (Sec. 4.2) across several quadruped categories. We demonstrate that our DualPM can be used to fit 3D skeletons for motion transfer and animations (Sec. 4.3), assess the impact of our design choices in an ablation study (Sec. 4.4), and showcase zero-shot generalization to unseen quadruped categories. Additionally, we discuss the limitations of our method and provide technical and implementation details in the supplementary material.
@@ 4.1 Qualitative evaluation
We qualitatively evaluate our method on challenging in-the-wild images of quadrupeds, specifically the horse, cow, sheep, and pig categories from the PASCAL VOC dataset [12] and goat images from [4]. The results are presented in Figs. 1 and 5. Our method successfully recovers the 3D shape and pose of quadrupeds from a single image, despite significant variations in pose, shape, and appearance within each animal category. For instance, in the horse category, although the method is trained on a single adult horse model, it generalizes well to foals and dwarf horses. We also compare our method with the state-of-the-art approach for single-view reconstruction of deformable objects, 3D-Fauna [33], and present side-by-side comparisons in Fig. 5. Our method performs better in recovering the shape and pose than 3D-Fauna, which struggles to accurately follow the object’s structure and fails to reconstruct intricate details specific to each object.
@@ 4.2 Quantitative evaluation
Following prior work on deformable object reconstruction [26, 30, 72, 33], we evaluate our method on the keypoint transfer task using the PASCAL VOC dataset [12]. We adhere to the protocol of [26] and report the percentage of correctly transferred keypoints with a 10% threshold (PCK@0.1) between pairs of images. The results are presented in Tab. 1, where we compare our method with state-of-the-art approaches.
We also measure the bi-directional Chamfer distance between the predicted point maps and the ground truth 3D shape on the Animodel dataset introduced in Farm3D [18]. This dataset comprises renderings of 3D models of horses, cows, and sheep in various poses. To ensure a fair comparison, we train a separate model on data that excludes the poses used in this benchmark. The performance of our method and others is reported in Tab. 1.
Our method consistently outperforms all other approaches, including self-supervised methods trained on large datasets of real images of quadruped categories (3D-Fauna [33]) and approaches trained on large-scale 3D asset datasets (Trellis [73]), demonstrating its ability to generalize to unseen poses and shapes. Moreover, it exhibits generalization to real-world images of the same category, despite being trained solely on synthetic data obtained from one or two templates per category. Additionally, it demonstrates some zero-shot generalization across categories, as shown in the supplementary material.
@@ 4.3 Skeleton fitting and animation
For each point $(\boldsymbol{p}_{i},\boldsymbol{q}_{i})$ of the DualPM $(P,Q)$, we identify the closest vertex of the mesh in the canonical space to the canonical point $\boldsymbol{q}_{i}$, then associate the point $(\boldsymbol{p}_{i},\boldsymbol{q}_{i})$ with the skinning weights of that vertex. For each bone $b\in B$, we identify subsets $\{(\boldsymbol{p}_{i},\boldsymbol{q}_{i})_{b}\}$, where the associated skinning weights between the bone $b$ and the points $(\boldsymbol{p}_{i},\boldsymbol{q}_{i})$ are higher than 0.5. We then compute the transformation that aligns the subset of canonical points $\{(\boldsymbol{q}_{i})_{b}\}$ to the subset of posed points $\{(\boldsymbol{p}_{i})_{b}\}$ using the Procrustes algorithm. This transformation is applied to the bone $b$ in the canonical space to derive its corresponding pose in the posed space. This process is repeated for all bones in the skeleton to fit the 3D skeleton to the posed points. We show examples of our fitted skeletons in Fig. 1.
@@ 4.4 Ablation study
Additionally, we investigate the effect of conditioning the predictions of the posed point maps $P$ on the input image features $F$ and the canonical point maps $Q$. When the model is conditioned on the image features $F$, it tends to overfit to the more variable features of the training dataset. In contrast, conditioning the posed point maps $P$ solely on the canonical point maps $Q$ leads to better generalization, particularly on out-of-distribution images, as we also demonstrate qualitatively in Fig. 7.
@@
Figure 7: Effect of different conditioning schemes. Conditioning solely on the canonical point maps $Q$, as opposed to using both the image features $F$ and the canonical point maps $Q$, results in better generalization, particularly in extreme out-of-distribution images.
@@
   $P$ conditioned on $F$ 3.70 ± $\pm\ 1.79$ 1.32 ± $\pm\ 0.61$
Table 2: Ablation study on Animodel Horses. We report the bi-directional Chamfer Distance in centimeters (lower is better $\downarrow$) for Animodel Horses. The evaluation follows the protocol as Tab. 1.
@@ 5 Conclusion
We presented DualPMs, a novel representation for monocular 3D shape and pose reconstruction of deformable objects. Predicting DualPMs is straightforward for a neural network, and they enable solving a variety of geometric tasks for deformable objects, including not only 3D shape and pose reconstruction but also 3D keypoint localization, skeleton fitting, and motion transfer. Despite being trained solely on synthetic data generated from one or two models per category, our method generalizes effectively to real images and outperforms state-of-the-art methods on the task of 3D shape and pose reconstruction of horses, cows, and sheep.
Acknowledgments.
We thank Paul Engstler for helpful discussions. Ben Kaye is supported by the EPSRC Centre for Doctoral Training in Autonomous Intelligent Machines and Systems, EP/S024050/1. Tomas Jakab and Andrea Vedaldi are supported by ERC-CoG UNION 101001212. Tomas Jakab is also supported by VisualAI EP/T028572/1.
@@ Acknowledgments.
Appendix A Generalization to unseen categories
Figure A8: Results on unseen categories. A version of our model trained only on the horse category also demonstrates robust generalization to the unseen categories such as cow and sheep, despite being trained solely with a single horse model.
Given the generalization capabilities of our method demonstrated within a single category, we analyze the generalization of a model trained on a single category to unseen categories. Specifically, we consider a model trained on horses and evaluate its performance on cow and sheep categories. We evaluate our approach using the same datasets as for the horses, PASCAL VOC [12] and Animodel [18], following the same evaluation protocol, with results reported in Tab. A3. Furthermore, we provide additional qualitative results on the same dataset in Fig. A8. Our method exhibits strong zero-shot generalization to these categories, outperforming state-of-the-art approaches on both datasets, despite being trained exclusively on a single horse model.
Method PCK (%) Chamfer Distance (cm)
Cow Sheep Real-Sized Normalized
Cow Sheep Cow Sheep
A-CSM [26] 26.3 28.6 6.71 ± $\pm\ 1.81$ 2.84 ± $\pm\ 0.77$ 2.35 ± $\pm\ 0.68$ 2.48 ± $\pm\ 0.70$
MagicPony [72] 42.5 41.2 7.22 ± $\pm\ 1.53$ 3.43 ± $\pm\ 0.73$ 2.53 ± $\pm\ 0.59$ 3.00 ± $\pm\ 0.68$
Farm3D [18] 40.2 36.1 6.91 ± $\pm\ 1.49$ 3.79 ± $\pm\ 0.55$ 2.41 ± $\pm\ 0.54$ 3.31 ± $\pm\ 0.49$
3D-Fauna [33] 9.19 ± $\pm\ 2.40$ 3.51 ± $\pm\ 0.88$ 3.20 ± $\pm\ 0.80$ 3.06 ± $\pm\ 0.76$
Ours 63.0 64.2 4.74 ± $\pm\ 1.40$ 2.32 ± $\pm\ 0.78$ 1.67 ± $\pm\ 0.55$ 2.03 ± $\pm\ 0.71$
Table A3: Evaluation on unseen cow and sheep categories. We evaluate on PASCAL VOS, reporting PCK@0.1 (higher is better $\uparrow$), and on Animodel [18], reporting the bi-directional Chamfer Distance in centimeters (lower is better $\downarrow$). Our model, trained solely data from a single horse model, outperforms state-of-the-art approaches, which were trained on data that included these specific categories.
Appendix B Limitations
Figure A9: Typical failure cases. We illustrate representative failure cases caused by (a) extreme viewpoints, (b) shapes and poses far from the training distribution, and (c-d) inaccuracies in the object segmentation masks.
Despite demonstrating surprising generalization, a current limitation of our method is that any additional synthetic 3D models added to the training dataset would have to be in the same canonical space as the training data. Addressing the challenging problem of aligning the canonical spaces of multiple 3D models would allows us to train on significantly larger datasets which could in turn lead to significant gains in performance for our proposed model. Another limitation of our method is that it is not specifically trained to handle occlusions caused by other objects. This is a limitation shared with other methods, such as 3D-Fauna [72, 33]. To address this, we plan to extend the data generation pipeline to include synthetic occlusions. Additionally, as the 3D reconstruction problem is often ambiguous for the unseen parts of objects, our method predicts only the expectation over all possible reconstructions, which can lead to unrealistic results for the invisible regions. We illustrate our typical failure cases in Fig. A9.
Appendix C Technical details